OpenStudy (firejay5):

I need help figuring this out of the geometric mean of this word problem. I'll draw the figure out. Medal will be rewarded for your patronage! :D

OpenStudy (firejay5):

 How far from Blare should the service station be located, so that the proposed road will perpendicular to the highway? How long will the new road be?

OpenStudy (firejay5):

@TuringTest

OpenStudy (anonymous):

\[\frac{ AD }{ AC }=\frac{ AC }{ AB }\] \[(AC)^{2}=AD*AB\] \[40^{2}=(50-x)50\] 1600=2500-50x 50x=2500-1600=900 x=900/50=18 service station is 18 miles from blare

OpenStudy (anonymous):

triangles are similar.